OpenStudy (anonomyous):

what is the center of the circle given by the equation (x-2)^2+(y+4)^2=6? Wouldnt i have to use the pythagorean thereom or am i wrong?

OpenStudy (anonomyous):

or the circle formula

OpenStudy (anonymous):

The circle formula: (x - h)^2 + (y - k)^2 = r^2 The center would be (h, k) with radius of r.

OpenStudy (anonymous):

So, here you have h = 2 and k = -4 and r = sqrt(6)

OpenStudy (anonymous):

All good now, @anonomyous ?

OpenStudy (anonomyous):

yes thank you

OpenStudy (anonymous):

Don't forget to hit "best response"! Thanks in advance! @anonomyous